student is using ANOVA to test 3 samples of data. Each sample has 6 data values. If she uses a - 0.05, what should she calculate as a critical value for her test? Select one: a. 3.5915 O b. 4.7571 OC.4.7650 d. 3.6823
Answer #1

You can find the critical values for an ANOVA hypothesis using the F-table. Because the F-distribution is based on two types of degrees of freedom, there’s one table for each possible value of alpha (the level of significance).

  • Numerator degrees of freedom = treatments – 1 = t – 1 = 3 – 1 = 2

  • Denominator degrees of freedom = total observations minus treatments = N – t = 3*6– 3 = 15


From the F Table, 2,15 = 3.6823